There are 6 routes from Delhi to Mumbai and 12 routes from Mumbai to Bangalore. In how many ways can you travel from Delhi to Bangalore via Mumbai?

18


72


36


16
1
Expert's answer
2020-10-01T09:56:17-0400

Given that,


Number of routes from Delhi to Mumbai a = 6

Number of routes from Mumbai to Bangalore b = 12


Hence the total number of routes from Delhi to Bangalore via Mumbai is given by c = aCr * bCr


Since one can travel in any 1 way only hence r = 1 in both the cases


"\therefore c = \frac {a!}{(a-r)!*r!} *\frac {b!}{(b-r)!*r!} = \frac {6!}{(6-1)!*1!}*\frac {12!}{(12-1)!*1!} = 72"


Hence we can travel in 72 different ways from Delhi to Bangalore via Mumbai.